# Jasmine: Harnessing Diffusion Prior for Self-Supervised Depth Estimation

### v1

@@ Appendix D Addition Experiments Results
Table 5: Ablation Studies on the hybird ratio. Ph refers to supervision using Eq. 7, while Latent refers to supervision using Eq. 6 (both in paper). The notation $+\lambda$ denotes the proportion of the KITTI dataset used (e.g., $+0.3$ indicates a KITTI:Hypersim ratio of 3:7 in the HIR training data.)
@@
In the paper, fig. 2 (g) illustrates the performance variation versus the hybrid ratio under two supervision schemes. We further present the complete results in Table 5. Note that although “Ph+0” offers better metrics, it predicts blurred results (7 (c-2)).
Qualitative Comparisons
@@
Appendix E Future Work
As noted in paper Sec. 2, the ubiquitously available videos suggest that the self-supervised methods possess significant data advantages and working potential. In this paper, Jasmine is trained on only tens of thousands of data samples, leaving room for further exploration in scaling up training data. We believe that if there really emerges a ‘GPT’ moment for 3D perception in the future, it will more likely involve self-supervised methods trained on the videos rather than learning from annotated depth.

### v2

@@ E.3 Qualitative Comparisons
Table 6: Ablation Studies on the mix-batch ratio. Ph refers to supervision using Eq. 5, while Latent refers to supervision using Eq. 4 . The notation $\lambda$ denotes the proportion of the KITTI dataset used (e.g., +0.3+0.3 indicates a KITTI:Hypersim ratio of 3:7 in the MIR training data.)
@@
E.4 Mix-batch Ratio Ablation Details
In the paper, Fig. 3 (g) illustrates the performance variation versus the mix-batch ratio under two supervision schemes. We further present the complete results in Table 6. Note that although “Ph+0” offers better metrics, it predicts blurred results (8 (c-2)).
Appendix F Analysis of SD Prior Degradation via Self-Supervision
For clarity, we first revisit the core definition of disparity: when capturing two images of the same scene from different camera positions, the same point will appear at different pixel coordinates in each image. This difference is known as disparity. In fact, disparity and depth are inversely proportional and correspond one-to-one.
From this perspective, the loss function in Eq. 1 is designed to find, for each point in the target view $I_{t}$, the most similar point in the source view $I_{t^{\prime}}$. From the coordinate difference between these points $I_{t^{\prime}\rightarrow t}$, we obtain the depth supervision (Eq. 2). However, if the corresponding point in $I_{t}$ is occluded in $I_{t^{\prime}}$, the optimization process is forced to select the "least bad" alternative, resulting in an incorrect match and, consequently, erroneous depth estimation.
[…]
Due to camera movement, the tree in the source view occludes the correct matching point of p. To minimize photometric loss, the algorithm searches for the most similar region nearby and once again finds the orange area at p’, which is 10 pixels (10m).
This error causes the expected depth edge on the right side of point p to disappear (right side’s depth are all 10m, reason same to p), resulting in a blurred boundary in the depth map. When such depth map is used as a supervisory signal to guide the SD model, it effectively introduces "noisy" data. This forces the SD model to learn and reproduce these incorrect and blurred boundaries, thereby undermining its strong prior knowledge of clear object boundaries.
Appendix G Evaluation Metrics
Similar to mdp2 , we employ the following evaluation metrics in our experiments,
[…]
RMSE: $\sqrt{\frac{1}{|{M}_{vl}|}\sum_{d\in{M}_{vl}}\left\|d-d_{gt}\right\|^{2}}$;
${a}_{t}$: percentage of $d$ such that $\max(\frac{d}{d_{gt}},\frac{d_{gt}}{d})<1.25^{t}$ ; where $d_{gt}$ and $d$ denote the GT and estimated pixel depth, ${M}_{vl}$ is the valid mask set to $1e-3<d_{gt}<80$.
Appendix H Error Bar Analysis
Figure 9: Error Bar Analysis on KITTI Eigen test split. We conduct this analysis through multiple training runs and observe the performance oscillations after 25k steps.
As highlighted in Sec 3.4, directly training Jasmine can be unstable due to the SD’s enormous size, joint training across modules, and indirect self-supervisory mechanisms. To further ensure stability, we also implement a module-wise freezing training strategy, which involves 3 key phases: In the beginning, we enable gradient updates for all network components. Subsequently, we freeze the SSG and Posenet modules to decouple depth-pose optimization while maintaining fixed parameters. After achieving convergence to a suboptimal solution, we reintroduce gradient updates to these modules for final optimization. In the training procedure, the first training phase involved 15k steps, while the second and third phases automatically transitioned, sharing an additional 10k steps for a total of 25k training steps. As illustrated in Fig. 9, we demonstrate that the pseudo-label supervision and module-wise freezing training are particularly crucial for steady training in complex, multi-module self-supervised systems.
Appendix I Limitation and Future Work
As noted in Sec. 2, the ubiquitously available videos suggest that the self-supervised methods possess significant data advantages and working potential. In this paper, Jasmine is trained on only tens of thousands of data samples and only on a driving dataset (KITTI), leaving room for further exploration in scaling up training data to other domains (industry, indoor, etc). We believe that if there really emerges a ‘GPT’ moment for 3D perception in the future, it will more likely involve self-supervised methods trained on videos rather than learning from annotated depth. Furthermore, we believe that the unsupervised Stable Diffusion fine-tuning paradigm proposed in this paper can be applied to other related fields, such as depth completion yan1 ; yan2 ; yan3 , depth super-resolution yan4 ; yan5 , and Multi-view Stereo yuan1 ; yuan2 ; yuan3 ; yuan4 .